# Neural-NPV Control: Learning Parameter-Dependent Controllers and Lyapunov Functions with Neural Networks

### 17 Mar 2026 v1

@@ IV-A Inverted pendulum with varying control effectiveness
where $x=[\phi,\dot{\phi}]^{T}$ and $u=\tau$ denote the state and control input (torque), respectively, $m=0.1$ $L=0.5$, $b=0.2$, and $g=9.81$ denote the mass, length, damping ratio, and gravitational constant. We impose an input constraint of $\mathinner{\!\left\lvert u(t)\right\rvert}\leq 3$ for all $t\geq 0$. The time-varying parameter $\theta(t)$ is introduced to characterize the control effectiveness of the actuator. A reduced $\theta(t)$ can represent actuator failure or other uncertain non-linearity of the system that affect the control effectiveness of the system. For instance, $\theta=0.3$ implies a loss of 70% control effectiveness compared to full control effectiveness corresponding to $\theta=1$. We assumed the following bounds: $\phi,\dot{\phi}]\in\mathcal{X}=[-\pi,\pi]\times[-6,6]$, $\theta(t)\in\Theta=[0.2,1]$ and $|\dot{\theta}(t)|\leq 0.1$. The structure of the Lyapunov and the controller network is shown in table Table I. Notice that the input to these networks are 4 channels rather than 3. In order to handle the discontinuity at $\phi=\pm\pi$ and provide a smooth, globally valid state representation for the network, we convert the angular state $\phi$ to $(\sin\phi,\,\cos\phi)$, which is fed to both control and Lyapunov networks together with $\dot{\phi}$ and $\theta$.
TABLE I: Network architectures and training hyperparameters.
@@
We compared Neural-NPV to the method in [30], referred to as SOS-NPV, which jointly synthesizes a PD controller and a PD Lyapunov function using SOS optimization. The method in [30] was slightly extended to incorporate input constraint following the idea in [29]. From Fig. 2, we can see that Neural-NPV achieves a robust ROA whose volume is markedly larger than that obtained by SOS-NPV, and covers the whole state space along the $\phi$ axis. In SOS-NPV, the Lyapunov function was parameterized as $V(x,\theta)=x^{T}X^{-1}(\phi,\theta)x$, where $X(\phi,\theta)$ is a polynomial matrix function. Notice that we intentionally made $X(\phi,\theta)$ not dependent on $\dot{\phi}$ (whose derivative is directly affected by $u$) to have a convex optimization problem [16, 13, 28, 20, 30]. In comparison, within Neural-NPV, the matrix term in the Lyapunov function 12 has a more general structure (not limited to the inverse of a polynomial function) and can depend on all states.
This demonstrates the superiority of our framework compared to SOS-based optimization methods. To demonstrate the performance of the learned controller, we used a time-varying disturbance trajectory defined by $\theta(t)=0.6+0.4\cos(t)$. The resulting state trajectories and Lyapunov function evolution are shown in Fig. 3. In our PGD-based verification experiments (with procedure detailed in Section III-C), we generated $10^{6}$ adversarial samples in $\Lambda^{\rho}_{V}\times\Theta$ and obtained a violation rate of approximately $1\%$ with a tolerance of $10^{-3}$, empirically validating $\Lambda^{\rho}_{V}$ as a robust ROA for the inverted pendulum system. Trajectory-based verification confirmed $100\%$ convergence to the equilibrium across all $10^{6}$ sampled initial conditions.
@@ IV-B Quadrotor with varying external disturbance
To demonstrate the scalability of Neural-NPV to high-dimensional systems, we apply it to a quadrotor system subject to external disturbances. The quadrotor model has the state-space representation, $x=\begin{bmatrix}p_{x},\ p_{y},\ p_{z},\ \dot{p}_{x},\ \dot{p}_{y},\ \dot{p}_{z}\end{bmatrix}^{\top}\in\mathbb{R}^{6}$, and three scheduling parameters $\theta(t)=\begin{bmatrix}\theta_{x},\theta_{y},\theta_{z}\end{bmatrix}\in\mathbb{R}^{3}$, denoting the external disturbance forces applied to the system. We assume that the disturbance is known. In practice, the disturbances can be estimated using an adaptive estimator [26] or a disturbance estimator [4]. The position & velocities are both expressed in the global inertial frame with a vertical axis pointing downward (North-East-Down convention). Adopting the NED frame for the quadrotor body, the attitude (roll, pitch, yaw) is parameterized as $(\phi,\theta_{p},\psi)$, and $\tau>0$ denotes the total thrust generated by the four rotors. The control input is defined as ${u}\triangleq\begin{bmatrix}\tau,\ \phi,\ \theta_{p}\end{bmatrix}^{\top}$. The outer-loop controller outputs desired thrust $\tau$ and attitude angles $(\phi,\theta_{p})$, which serve as references for an inner-loop attitude controller assumed to operate on a much faster timescale. As yaw is not actively controlled; we set $\psi=0$. Given this parameterization, the translational dynamics of the quadrotor subject to the external disturbance $\theta(t)=[\theta_{x},\theta_{y},\theta_{z}]^{\top}\in\mathbb{R}^{3}$ are
$\begin{bmatrix}\ddot{p}_{x}\\ \ddot{p}_{y}\\ \ddot{p}_{z}\end{bmatrix}=g\mathbf{e}_{3}-\frac{\tau}{m}Re_{3}+\frac{\theta}{m}=\begin{bmatrix}-\frac{\tau}{m}\cos\phi\sin\theta_{p}+\frac{\theta_{x}}{m}\\ \frac{\tau}{m}\sin\phi+\frac{\theta_{y}}{m}\\ g-\frac{\tau}{m}\cos\theta_{p}\cos\phi+\frac{\theta_{z}}{m}\end{bmatrix}$, (22)
@@
where $g$ is the local gravitational acceleration, $\mathbf{e}_{3}=[0,0,1]^{\top}$, and $R$ is the rotation matrix that transforms thrust from body frame to the inertial frame following a ZYX (yaw-pitch-roll) rotation sequence. The following state and input constraints are imposed to ensure physical feasibility and permit sufficiently aggressive maneuvers: $(\phi,\,\theta_{p})\in[-90^{\circ},\,90^{\circ}]\times[-180^{\circ},\,180^{\circ}],\tau\in[0\,,\,11.3]\text{N}$. From 22, we can see that the dynamics are not affine in control. This prevents the direct application of SOS-based methods which are only applicable to control-affine systems [10, 12, 30].
@@
We consider the following bounds: $p_{x},p_{y},p_{z},\dot{p}_{x},\dot{p}_{y},\dot{p}_{z}]\in\mathcal{X}=[-6,6]^{6},\theta(t)\in\Theta=[-2,2]^{3},\ \dot{\theta}(t)\in\Omega=[-0.5,0.5]^{3}$. The details of the Lyapunov network and the controller network can be found in Table I. To test the performance of the controller, we used a time-varying disturbance trajectory defined by $\theta(t)=\begin{bmatrix}2\sin(\pi t),2\cos(\frac{3}{4}\pi t),2\cos(\frac{\pi}{2}t)\end{bmatrix}^{T}$, which simulates a realistic external disturbance acting on the quadrotor in the inertial frame. The resulting state trajectories and Lyapunov function evolution are shown in Fig. 5. We can observe in Fig. 6 that the control inputs are non-zero even when the states have already converged to the equilibrium in Fig. 5, showcasing the learned PD controller’s capability of handling the external disturbances robustly.
@@
To verify our robust ROA empirically via the procedure described in Section III-C, we sampled $10^{6}$ trajectories in $\Lambda^{\rho}_{V}\times\Theta$. The controller converged all sampled trajectories to the equilibrium, demonstrating the empirical robustness of the synthesized controller. For PGD-based verification, we generated $10^{6}$ adversarial samples in $\Lambda^{\rho}_{V}\times\Theta$ and find that the counterexample violation rate obtained is approximately $1\%$ across all evaluation rounds with a tolerance of $10^{-3}$. Both of our empirical verification schemes validate $\Lambda^{\rho}_{V}$ as a robust ROA for the quadrotor system.
@@
TABLE II: Empirical verification results under two evaluation schemes.
System Scheme Samples Violation rate Convergence (%)
Inv. Pendulum PGD $0^{6} 11$
Trajectory $10^{6}$ 100
Quadrotor PGD $0^{6} 11$
Trajectory $10^{6}$ 100
@@ V Conclusions
In this paper, we presented Neural-NPV, a novel learning-based framework to jointly synthesize and verify a parameter-dependent (PD) Lyapunov function and PD controller for nonlinear parameter-varying (NPV) systems. Our two-stage training framework uses a counterexample-guided training procedure, and promotes the growth of the certified robust region of attraction. The framework is applicable to control-non-affine systems,is scalable to high-dimensional and handles input constraints seamlessly. . It is validated on two platforms, and simulation results as well as our empirical verification schemes demonstrate its advantages over SOS-based methods. One limitation of our current approach is the lack of formal verification of the learned systems through neural network verifiers such as $\alpha,\beta$-CROWN. For future work, we intend to formally verify the learned controller and Lyapunov functions, and experimentally validate the proposed approach on a real quadrotor system.

### 07 Apr 2026 v2

@@ IV-A Inverted pendulum with varying control effectiveness
where $x=[\phi,\dot{\phi}]^{T}$ and $u=\tau$ denote the state and control input (torque), respectively, $m=0.1$ $L=0.5$, $b=0.2$, and $g=9.81$ denote the mass, length, damping ratio, and gravitational constant. We impose an input constraint of $\mathinner{\!\left\lvert u(t)\right\rvert}\leq 3$ for all $t\geq 0$. The parameter $\theta(t)$ is introduced to characterize the control effectiveness of the actuator.
We assumed the following bounds: $\phi,\dot{\phi}]\in\mathcal{X}=[-\pi,\pi]\times[-6,6]$, $\theta(t)\in\Theta=[0.2,1]$ and $|\dot{\theta}(t)|\leq 0.1$. The structure of the Lyapunov and the controller network is shown in table Table I. Notice that the input to these networks are 4 channels rather than 3. In order to handle the discontinuity at $\phi=\pm\pi$ and provide a smooth, globally valid state representation for the network, we convert the angular state $\phi$ to $(\sin\phi,\,\cos\phi)$, which is fed to both control and Lyapunov networks together with $\dot{\phi}$ and $\theta$.
TABLE I: Network architectures and hyperparameters
@@
We compared Neural-NPV to SOS-NPV [26], which jointly synthesizes a PD controller and PD Lyapunov function via SOS optimization, extended here with input constraints following [25]. From Fig. 2, we can see that Neural-NPV achieves a robust ROA whose volume is markedly larger than that obtained by SOS-NPV, and covers the whole state space along the $\phi$ axis. In SOS-NPV, the Lyapunov function was parameterized as $V(x,\theta)=x^{T}X^{-1}(\phi,\theta)x$, where $X(\phi,\theta)$ is a polynomial matrix function. Notice that we intentionally made $X(\phi,\theta)$ not dependent on $\dot{\phi}$ (whose derivative is directly affected by $u$) to have a convex optimization problem [14, 26]. In comparison, within Neural-NPV, the matrix term in the Lyapunov function 12 has a more general structure (not limited to the inverse of a polynomial function) and can depend on all states. This demonstrates the superiority of our framework compared to SOS-based optimization methods. To demonstrate the performance of the learned controller, we used a time-varying disturbance trajectory defined by $\theta(t)=0.6+0.4\cos(t)$. The resulting state trajectories and Lyapunov function evolutions are shown in Fig. 3 Trajectory-based verification confirmed $100\%$ convergence to the equilibrium across all $10^{6}$ sampled initial conditions. In our PGD-based verification experiments (with procedure detailed in Section III-C), we generated $10^{6}$ adversarial samples in $\Lambda^{\rho}_{V}\times\Theta$ and obtained a violation rate of approximately $1\%$ with a tolerance of $10^{-3}$, empirically validating $\Lambda^{\rho}_{V}$ as a robust ROA for the inverted pendulum system.
@@ IV-B Quadrotor with varying external disturbance
To demonstrate the scalability of Neural-NPV to high-dimensional systems, we apply it to a quadrotor system subject to external disturbances:
[…]
@@
where the state vector is $x=\begin{bmatrix}p_{x},\ p_{y},\ p_{z},\ \dot{p}_{x},\ \dot{p}_{y},\ \dot{p}_{z}\end{bmatrix}^{\top}\in\mathbb{R}^{6}$, and the scheduling parameter vector $\theta(t)=\begin{bmatrix}\theta_{x},\theta_{y},\theta_{z}\end{bmatrix}\in\mathbb{R}^{3}$ denotes the external disturbance forces along $x$, $y$, and $z$ axes. Here, we assume that the disturbance is known. In practice, the disturbances can be estimated using an adaptive estimator [22] or a disturbance estimator [4]. The attitude (roll, pitch, yaw) is parameterized as $(\phi,\theta_{p},\psi)$, and $\tau>0$ denotes the total thrust generated by the four rotors. The control input is defined as ${u}\triangleq\begin{bmatrix}\tau,\ \phi,\ \theta_{p}\end{bmatrix}^{\top}$. The outer-loop controller outputs desired thrust $\tau$ and attitude angles $(\phi,\theta_{p})$, which serve as references for an inner-loop attitude controller assumed to operate on a much faster timescale. As yaw is not actively controlled, we set $\psi=0$.
@@
In 22, $g=9.81m/s^{2}$, $\mathbf{e}_{3}=[0,0,1]^{\top}$, and $R$ is the rotation matrix from body frame to the inertial frame following a ZYX (yaw-pitch-roll) sequence. The following state and input constraints are imposed: $(\phi,\,\theta_{p})\in[-90^{\circ},\,90^{\circ}]\times[-180^{\circ},\,180^{\circ}],\tau\in[0\,,\,11.3]\text{N}$. From 22, we can see that the dynamics are not affine in control. This prevents the direct application of SOS-based methods which are only applicable to control-affine systems [8, 10, 26].
@@
We consider the following ROIs: $x\in\mathcal{X}=[-6,6]^{6},\theta(t)\in\Theta=[-2,2]^{3},\ \dot{\theta}(t)\in\Omega=[-0.5,0.5]^{3}$. The details of the Lyapunov network and the controller network can be found in Table I. To test the performance of the controller, we used a time-varying disturbance trajectory defined by $\theta(t)=\begin{bmatrix}2\sin(\pi t),2\cos(\frac{3}{4}\pi t),2\cos(\frac{\pi}{2}t)\end{bmatrix}^{T}$, which simulates an external disturbance acting on the quadrotor in the inertial frame. The resulting state trajectories and Lyapunov function evolution are shown in Fig. 5. From Fig. 6, we can see that after the states have already converged to the equilibrium in Fig. 5, the controller still constantly adjusts the control inputs to compensate for the external disturbances in order to keep the system at the equlibrium.
@@
To verify our robust ROA empirically via the procedure described in Section III-C, we sampled $10^{6}$ trajectories in $\Lambda^{\rho}_{V}\times\Theta$. The controller drove the system to the equilibrium under all cases, empirically demonstrating the robustness of the controller. For PGD-based verification, we generated $10^{6}$ adversarial samples in $\Lambda^{\rho}_{V}\times\Theta$ and find that the counterexample violation rate obtained is approximately $1\%$ across all evaluation rounds with a tolerance of $10^{-3}$. Both of our empirical verification schemes validate $\Lambda^{\rho}_{V}$ as a robust ROA for the quadrotor system.
@@
TABLE II: Empirical verification results under two evaluation schemes described in Section III-C
System Scheme Samples Violation rate Convergence
Inv. Pendulum PGD $0^{6} 1~1\%$
Trajectory $10^{6}$ 100%
Quadrotor PGD $0^{6} 11\%$
Trajectory $10^{6}$ 100%
@@ V Conclusions
We presented Neural-NPV, a novel learning-based framework to jointly synthesize and verify a parameter-dependent (PD) Lyapunov function and PD controller for nonlinear parameter-varying (NPV) systems. Our two-stage training framework uses a counterexample-guided training procedure, and promotes the growth of the certified robust region of attraction. The framework is applicable to control-non-affine systems, is scalable to high-dimensional systems, and handles input constraints seamlessly. Numerical experiments on an inverted pendulum and a quadrotor system demonstrate the effectiveness of our proposed approach. One limitation of our current approach is the lack of formal verification of the learned systems. For future work, we intend to formally verify the learned controller and Lyapunov functions using neural network verifiers and experimentally validate the proposed approach on a real quadrotor system.